I am graphing quadratic functions. My equation is f(x)= x^2+5. What do I do?
Answer by Edwin McCravy(20054)

Make a table of values and get 7 point:

[Note that y and f(x) mean the same thing]

 x |     x²+5      y   point
-------------------------------------
-3 |(-3)²+5=9+5 = 14  (-3,14)
-2 |(-2)²+5=4+5 =  9   (-2,9)
-1 |(-1)²+5=1+5 =  6   (-1,6)
 0 | (0)²+5=0+5 =  5    (0,5)
 1 | (1)²+5=1+5 =  6    (1,6)
 2 | (2)²+5=4+5 =  9    (2,9)
 3 | (3)²+5=9+5 = 14   (3,14)

Plot them:



Draw a U-shaped graph through them:



Edwin